OpenStudy (anonymous):

Put the following radical expression into simplified form. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\sqrt{\frac{ 36a^4 }{ 7 }}\]

OpenStudy (anonymous):

I know 36=6 & a^4=a^2, but the square root of a^2 is a, sooooo?

hartnn (hartnn):

so your numerator is just 6a^2 right ?? how will you eliminate square root in denominator ? you've done this type of problem before...

OpenStudy (anonymous):

multiply it by 6a^2

OpenStudy (anonymous):

To eliminate the the square root of the denominator? :P

OpenStudy (anonymous):

Oooh no, by 7, sorry

hartnn (hartnn):



OpenStudy (anonymous):

^^

OpenStudy (anonymous):

42a/7

OpenStudy (anonymous):

square root of 7

hartnn (hartnn):

there's no 42  thats it!